# Reliable Fine-Grained Evaluation of Natural Language Math Proofs

### v1

@@ 3.5 In-depth Analysis
Comparing evaluators based on o3 with None to those given a Ref+MS, we find that in over $60\%$ of cases, the no-context evaluator overestimates the generation’s correctness, i.e., it incorrectly assigns a higher score to the generation. Crucially, this overestimation is not uniform: we observe a strong correlation ($r=0.699$, $p<0.001$) between proof quality and evaluation gap. For low-quality proofs (scored 0–2 by the context-aware evaluator), the no-context evaluator over-scores by an average of 1.7 points. In stark contrast, for high-quality proofs (scored 5–7), the context-aware evaluator assigns higher scores by +0.8 points on average. This asymmetric pattern supports our hypothesis that, without reference solutions or marking schemes, the evaluator struggles to gauge proof progress—how much has been correctly established and how far the argument is from completion—and thus is prone to rewarding fluent but irrelevant or unhelpful content.
This effect is exacerbated when the evaluator cannot fully solve the problem—a case that constitutes the majority of our dataset, as shown in Figure 3. When we examine problems that o3 (the evaluator model) cannot solve well (scoring $\leq 2$ when acting as a prover), we find that the no-context evaluator over-scores by 1.40 points on average. In contrast, for problems that o3 can solve (scoring $\geq 5$), the context-aware evaluator actually scores higher by 0.90 points. This 2.3-point swing (t=8.790t=-8.790, $p<0.001$) demonstrates that the evaluator’s difficulty in solving a problem directly impairs its ability to assess partial progress on that problem without external guidance.
@@
In our main setup, evaluators use the same marking schemes as the human graders. To test sensitivity to the rubric, we replace these with two alternatives: (i) regenerate a marking scheme using the same prompt/model (new sample), and (ii) generate a marking scheme with a different model. We then evaluate the dataset with o3 given the reference solution plus each alternative scheme. In both cases, performance degrades relative to using the original marking scheme across different metrics, suggesting that evaluator accuracy depends on close alignment with the marking scheme used by human although experts do not strictly follow them. Detailed results appear in §A.4.
@@ 4 Downstream Utility: Best-of-N Proof Selection
A key property of a good reward model is that it can reliably identify the best response in a given batch. To understand how this ability correlates with the calibration metrics in Section 3.2, we measure the best-of-$n$ (BoN) performance of various evaluators. BoN is conditioned on the generator model and the sampling budget ($n$), directly reflecting the operations that drive model improvement cycles. It can reveal issues like over-optimization to a flawed evaluator and yields a practical utility curve as $n$ varies. In general, a higher BoN score means the evaluator more reliably selects high-quality candidates for subsequent training, guiding real downstream gains. The applications include selecting high-quality training data via rejection sampling (e.g., distillation)  (Nakano et al., 2022; Touvron et al., 2023) and providing a reward signal for Reinforcement Learning (RL) (DeepSeek-AI et al., 2025; Yu et al., 2025).
(a) Evaluators as BoN selectors. Ensemble-based fine-grained evaluator closely track the Human-Oracle curve, while the binary evaluators (green lines) perform much worse.
@@
To create a testbed, we use o3 to generate 16 candidate proofs for each of 29 selected problems from 2025, resulting in 464 unique proofs. All 464 of these candidates were then scored by three human experts using the pipeline described in §2. This dataset is distinct from the section 3 split and serves a different purpose. For each problem, 8 responses are graded by two experts with disagreements resolved through discussion. For this analysis, we test a selection of the fine-grained evaluators studied previously (§4.1) and several comparison-based selection strategies including Knockout tournament style selection (§4.2). The performance of each selection method is measured using a best-of-$n$ (BoN) curve. Since an exhaustive evaluation over all $\binom{16}{n}$ subsets is computationally intensive, we estimate the BoN curve using Monte Carlo subsampling. For each $n\in\{1,2,\cdots,16\}$, we sample a large number of subsets of size $n$ without replacement. For each subset, the evaluator selects the single proof with the highest assigned score, using the initial ordering to break ties. The performance at $n$ is the average human score of these selected proofs, aggregated over all subsamples and all 29 problems. We also include two key baselines: an oracle baseline and a mean baseline. The Human Oracle represents the performance ceiling, calculated as the average score of the best possible selection among the first $n$ candidates. The Mean Baseline is calculated as the cumulative average score of the first $n$ candidates. All evaluators and selectors use o3 as the backbone model.
@@ 4.1 The Value of Fine-Grained Scoring
Results are shown in 4(a), addressing three key questions.
How close is ProofGrader to the human oracle?
Our ProofGrader (an ensemble of five o3 runs with Ref+MS) closely tracks the human-oracle curve, consistently selecting higher-quality proofs as $n$ increases. At nn=16, it achieves an average score of 4.14.
Is the ranking between evaluators from §3 preserved?
Comparing the four different fine-grained evaluators from §3: Ref+MS ensemble (ProofGrader), Ref+MS, Ref, and None, we find the same ranking holds. Notably, None performs significantly worse than those using a marking scheme, confirming that the marking scheme is a critical component.
Do fine-grained scores provide a stronger selection signal than binary labels?
We further compare with binary evaluators: the o3-based models prompted to classify each proof as “Correct” or “Incorrect” rather than assigning a 0–7 score.
In contrast to ProofGrader, which reaches a score of 4.14 at nn=16, the binary evaluator performs only slightly better than the average score of all candidates. This limitation comes from collapsing all “correct” (or “incorrect”) proofs into a single category, losing the ability to rank solutions. When multiple correct candidates exist, it cannot distinguish an adequate proof (e.g., a 5/7) from an excellent one (7/7). We thus conclude that fine-grained scoring preserves relative ordering, making it essential for effective reward models in mathematical reasoning.
@@ 4.2 Robustness Against Comparison-based Selection Strategies
Tournament. For candidates $s_{1},\dots,s_{n}$}, the selector performs pairwise comparisons for every unordered pair $(i,j)$. Let $W_{ij}\in{0,1}$ indicate whether $s_{i}$ beats $s_{j}$. We form a win-rate matrix ($W$) and compute $w_{i}=\sum_{j\neq i}W_{ij}$. The selected answer is $\arg\max_{i}w_{i}$ (tie-break by head-to-head or margin if applicable). This uses $\binom{n}{2}$ comparisons.
Knockout. Candidates are placed in a bracket; the selector compares pairs, and the winner advances until one remains. This requires (n1n-1) comparisons per bracket. (Optionally, we average over multiple random brackets and select the modal winner or the highest aggregate win rate.)
@@ 5 Related Work
While prior work has established the importance of proof-based evaluation and explored the “LLM-as-a-judge” paradigm, a critical gap remains. No existing work has conducted a systematic, empirical study of the evaluator design space for mathematical proofs, nor provided a comprehensive dataset with the fine-grained, expert-annotated proofs necessary for such a study. Consequently, the key factors that determine evaluator robustness remain largely uninvestigated. Our work addresses these gaps, and we situate our contributions in relation to several primary lines of research below.
@@
Datasets such as GSM8K (Cobbe et al., 2021), which targets grade-school word problems, and MATH (Hendrycks et al., 2021) and AIME, which extend coverage to algebra, calculus, and contest problems, have served as important benchmarks in math reasoning, but primarily focus on short, closed-form answers. More recent competition based benchmarks, including Omni-MATH (Gao et al., 2024), OlympiadBench (He et al., 2024), HARP (Yue et al., 2024), and MathArena (Balunović et al., 2025), are substantially more challenging; however, they still largely emphasize final answer matching, leaving mathematical proof problems underrepresented. MathArena includes proof problems, but only focuses on latest competitions. Meanwhile, proof generation has advanced rapidly—Gemini (Luong & Lockhart, 2025) and unreleased OpenAI systems have reached IMO gold level performance, and model-agnostic verification-and-refinement enables comparable results for base models (Huang & Yang, 2025). However, these studies largely depend on manual grading to assess proof quality (Petrov et al., 2025).
@@
LLM-as-a-judge methods score open-ended outputs along axes such as correctness, enabling scalable evaluation at reduced human cost (Zheng et al., 2023; Li et al., 2024). Subsequent work probes judge reliability on harder tasks (Tan et al., 2025; Frick et al., 2024) and improves fidelity with prompt-specific rubrics that encode task-adapted criteria (e.g., relevance, depth) directly in the prompt (Kim et al., 2024; Liu et al., 2023b; Wadhwa et al., 2025). For mathematical proofs, automated evaluation remains limited: prior efforts are either domain-specific (e.g., inequality proofs in IneqMath (Sheng et al., 2025)) or provide only coarse labels (binary correctness in Dekoninck et al. (2025)).
@@
A complementary line of work leverages interactive theorem provers (ITPs), where LLMs generate proofs in formal languages like Lean or Isabelle. The correctness of these proofs can be automatically verified by the ITP, guaranteeing their logical soundness. Major benchmarks in this area include miniF2F (Zheng et al., 2021), FIMO (Liu et al., 2023a), PutnamBench (Tsoukalas et al., 2024), which formalize competition math problems; and LeanWorkbook (Ying et al., 2024), which provides a broad Lean corpus for training and evaluation. While this approach is promising, our work focuses on informal proofs for two reasons. First, natural language remains the primary medium for mathematical communication in education and research. Second, the task of automatically translating informal proofs into a formal language, i.e., autoformalization, is itself an exceptionally challenging research problem (Gao et al., 2025; Liu et al., 2025a).
Evaluation for Reward Modeling.
Evaluating reward models (RMs) is crucial for ensuring their alignment with human preferences in downstream tasks. Early works pioneered the use of best-of-$n$ (BoN) sampling as a lightweight proxy for RM quality, where multiple LM generations are scored and the top-$n$ selected, with performance measured via win rates against gold-standard RMs (Stiennon et al., 2022; Nakano et al., 2022). Subsequent studies explored BoN’s vulnerabilities, such as overoptimization and reward hacking, through scaling experiments on preference datasets like HH-RLHF (Gao et al., 2022). Domain-specific evaluations, such as in math reasoning, leverage pairwise RMs for BoN tournaments, boosting accuracies on benchmarks, such as MATH-500 (Liu et al., 2025b). Comprehensive benchmarks like RewardBench 2 further correlate BoN performance with RM scores on diverse tasks (Malik et al., 2025).
@@ 6 Conclusion
In this work, we addressed the critical challenge of reliably evaluating natural language mathematical proofs. We introduce a systematic methodology for developing and validating fine-grained evaluators and create ProofBench, the first comprehensive annotated dataset of fine-grained proof grading that spans multiple contests and years. Using ProofBench, we systematically explore the evaluator design space and find that quality depends on a strong model backbone, informative context, and problem-specific marking schemes, and that ensembling further improves accuracy and robustness. Our analysis yields ProofGrader, the strongest evaluator considering all design choices. Finally, we demonstrate the practical utility of the evaluators in a downstream best-of-$n$ selection task, where it closes 78% of the performance gap to the human oracle. We expect our methodology, benchmark, and findings to provide a strong foundation for future research in hard-to-verify mathematical reasoning tasks.
Limitation and Future Work.
@@
While our work advances reliable evaluation of natural-language mathematical proofs, several limitations remain. First, our scope—olympiad-style proofs in natural language—does not yet cover research-grade arguments, specialized domains (e.g., algebraic geometry), or applied settings; extending to these regimes is an important next step. Second, evaluator quality can likely be improved through better prompting and workflow designs, Our dataset and analyses are intended as a foundation for such work, and we encourage future work to continue improve evaluators using ProofBench as a testbed. Our evaluation focuses exclusively on mathematical correctness. Other crucial aspects of natural-language proofs, such as readability, clarity, and elegance, are not currently assessed, and we leave the development of metrics for these qualities as future work. Third, our strongest evaluators currently rely on closed models. Although our design is model-agnostic, further improving fully open-source evaluators to match those based on closed models is an important direction. To support this, we release prompts and evaluation scripts for reproduction with open models. Beyond evaluation, our graders could be used to synthesize supervision for reward-model training. While our best-of-$n$ experiments use the evaluator as a selection signal, a first step toward training-time integration, we have not yet train with it; incorporating it into proof generator training is a promising future direction.
Appendix A Appendix
A.1 Data Information
@@
Contest Description
APMO The Asia Pacific Mathematical Olympiad is a regional proof-based contest for high school students across Asia–Pacific. It features five challenging problems solved over a fixed time window, proctored locally.
EGMO The European Girls’ Mathematical Olympiad is an international two-day proof contest for female students, modeled after the IMO. It promotes participation and excellence among young women in mathematics.
IMO The International Mathematical Olympiad is the premier global high school math competition. Over two days, students tackle six proof problems that test creativity, rigor, and depth.
PUTNAM The William Lowell Putnam Mathematical Competition is a highly challenging proof exam for undergraduates in the U.S. and Canada. It consists of 12 problems emphasizing ingenuity and precise argumentation.
TST Team Selection Tests are national-level olympiad exams used by USA to select their IMO teams. Problems mirror international difficulty and assess readiness for global competition.
USAMO The United States of America Mathematical Olympiad is an invitational proof contest following AMC/AIME qualification. It identifies top U.S. high school problem solvers and feeds into further training and team selection.
@@
Table 8: Core problem counts by contest and year.
A.2 Annotation Pipeline Details
Stage 1: Marking Scheme Finalization.
Pilot. We compare marking scheme generators from two model families, each with and without in-context examples, using an initial prompt distilled from authoritative grading materials.
Quality rating. For each problem–solution pair, annotators independently rate the generated marking scheme on a 0–3 scale (0 = invalid; 3 = high-fidelity), then discuss to consensus.
Selection and refinement. We select the best configuration (gemini-2.5-pro, no in-context example), iteratively refine the prompt based on annotator feedback, and re-evaluate on additional problems until agreement stabilizes. We then freeze the marking scheme generator for subsequent use.
Stage 2: Solution annotation.

### v2

@@ 3.5 In-depth Analysis
What are the common failure modes of ProofGrader?
To investigate failure cases, we manually inspected 50 solutions where 2 out of 5 ensemble runs deviated from human scores by at least 2 points. We find that over-crediting (10.8%) and under-crediting (12.2%) occur at similar rates. Over-scoring typically arises when ProofGrader relies on surface-level heuristics: it rewards solutions that mirror rubric structure or employ sophisticated frameworks (e.g., custom transformations) even when central claims are false. Furthermore, it often misclassifies fatal logical errors, such as broken reductions, as minor justification issues. Under-scoring commonly results from penalizing early missteps in otherwise correct proofs, double-counting a single flaw, or demanding excessive micro-justifications. By domain, failures are most frequent in Algebra and Geometry ($\sim$25% each). See §H for details.
@@
Comparing evaluators based on o3 with None to those given a Ref+MS, we find that in over $60\%$ of cases, the evaluator without context overestimates the generation’s correctness, i.e., it incorrectly assigns a higher score to the generation. Crucially, this overestimation is not uniform: we observe a strong correlation ($r=0.699$, $p<0.001$) between proof quality and evaluation gap. For low-quality proofs (scored 0–2), the no-context evaluator over-scores by an average of 1.7 points. In contrast, for high-quality proofs (scored 5–7), the evaluator with context information assigns higher scores by 0.8 points on average. This asymmetric pattern supports our hypothesis that, without reference solutions or marking schemes, the evaluator struggles to gauge proof progress.
@@
In our main setup, evaluators use the same marking schemes as the human graders. To test sensitivity to the rubric, we replace these with two alternatives: (i) regenerate a marking scheme using the same prompt/model (new sample), and (ii) generate a marking scheme with a different model. We then evaluate the dataset with o3 given the reference solution plus each alternative scheme. In both cases, performance degrades relative to using the original marking scheme across different metrics, suggesting that evaluator accuracy depends on close alignment with the marking scheme used by human although experts do not strictly follow them. Detailed results appear in §H.
@@ 4 Downstream Utility: Best-of-N Proof Selection
(a) Evaluators as BoN selectors. Ensemble-based fine-grained evaluator closely tracks the Human-Oracle curve, while the binary evaluators (green lines) perform much worse.
@@
A good reward model should be able to reliably identify the best response in a given batch. To understand how this ability correlates with the calibration metrics in Section 3.2, we measure the best-of-$n$ (BoN) performance of various evaluators. In general, a higher BoN score means the evaluator more reliably selects high-quality candidates for subsequent training, guiding real downstream gains. The applications include selecting high-quality training data via rejection sampling (e.g., distillation)  (Nakano et al., 2022; Touvron et al., 2023) and providing a reward signal for Reinforcement Learning (RL) (DeepSeek-AI et al., 2025; Yu et al., 2025).
@@
To create a testbed, we use o3 to generate 16 candidate proofs for each of 29 selected problems from 2025, resulting in 464 unique proofs. All 464 of these candidates were then scored by three human experts using the pipeline described in §2. This dataset is distinct from ProofBench and serves a different purpose. For each problem, 8 responses are graded by two experts with disagreements resolved through discussion. We test a selection of the fine-grained evaluators studied previously (§4.1) and several comparison-based selection strategies (§4.2). The performance of each selection method is measured using a best-of-$n$ (BoN) curve. Since an exhaustive evaluation over all $\binom{16}{n}$ subsets is computationally intensive, we estimate the BoN curve using Monte Carlo subsampling. For each $n$, we sample a large number of subsets of size $n$ without replacement. For each subset, the evaluator selects the single proof with the highest assigned score. The performance at $n$ is the average human score of these selected proofs, aggregated over all subsamples and all 29 problems. We also include two key baselines: an oracle baseline and a mean baseline. The Human Oracle represents the performance ceiling, calculated as the average score of the best possible selection among the sampled $n$ candidates. The Mean Baseline is calculated as the cumulative average score of the sampled candidates. All evaluators and selectors use o3 as the backbone model.
@@ 4.1 The Value of Fine-Grained Scoring
As shown in 4(a), our ProofGrader (an ensemble of five o3 runs with Ref+MS) closely tracks the human-oracle curve, consistently selecting higher-quality proofs as $n$ increases. At nn=16, it achieves an average score of 4.14. Comparing the four different fine-grained evaluators from §3: Ref+MS ensemble (ProofGrader), Ref+MS, Ref, and None, we find the same ranking holds. Notably, None performs significantly worse than those using a marking scheme, confirming that the marking scheme is a critical component.
We further compare with binary evaluators: the o3-based models prompted to classify each proof as “Correct” or “Incorrect” rather than assigning a 0–7 score. In contrast to fine-grained evaluators, the binary evaluators perform much worse. This limitation comes from collapsing all “correct” (or “incorrect”) proofs into a single category, losing the ability to rank solutions. When multiple correct candidates exist, it cannot distinguish an adequate proof (e.g., a 5/7) from an excellent one (7/7). We thus conclude that fine-grained scoring preserves relative ordering, making it essential for effective reward models in mathematical reasoning.
@@ 4.2 Robustness Against Comparison-based Selection Strategies
Tournament. For candidates $s_{1},\dots,s_{n}$}, the selector performs pairwise comparisons for every unordered pair $(i,j)$. Let $W_{ij}\in{0,1}$ indicate whether $s_{i}$ beats $s_{j}$. We form a win-rate matrix $W$ and compute $w_{i}=\sum_{j\neq i}W_{ij}$. The selected answer is $\arg\max_{i}w_{i}$ (tie-break by head-to-head or margin if applicable). This uses $\binom{n}{2}$ comparisons.
Knockout. Candidates are placed in a bracket; the selector compares pairs, and the winner advances until one remains. This requires n1n-1 comparisons per bracket. Similar to other approaches, for each $n$, we average the scores over multiple random brackets.
@@ 5 Related Work
GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and AIME provide math reasoning benchmarks but mainly target short, closed-form answers. More recent competition-based benchmarks: Omni-MATH (Gao et al., 2024), OlympiadBench (He et al., 2024), HARP (Yue et al., 2024), and MathArena (balunović2025matharenaevaluatingllmsuncontaminated) are more challenging yet still emphasize answer matching, leaving proof problems underrepresented. MathArena includes proofs but focuses only on latest competitions. Meanwhile, proof generation has advanced rapidly: Gemini (Luong and Lockhart, 2025), unreleased OpenAI systems, and model-agnostic verify-and-refine methods achieve IMO gold-level performance even for base models (Huang and Yang, 2025), but evaluations still mostly rely on manual grading (Petrov et al., 2025).
@@
LLM-as-a-judge methods use LLMs to score open-ended responses along dimensions such as correctness, enabling scalable evaluation while reducing human annotation costs (Zheng et al., 2023a; Li et al., 2024). Recent work has further examined the reliability of LLM judges on more challenging tasks (Tan et al., 2025; Frick et al., 2024). For mathematical proofs, however, automated evaluation remains relatively underexplored. Earlier work (Welleck et al., 2022; Frieder et al., 2023) often relies on human evaluation or lexical automatic metrics, which are difficult to scale and insufficient for verifying complex, competition-level proofs. Existing approaches are also often highly specialized, such as IneqMath (Sheng et al., 2025) for inequality proofs, or lack fine-grained evaluation; for example, the dataset of Dekoninck et al. (2025) provides only binary annotations. As a result, many studies still depend on manual evaluation for model-generated proofs (Petrov et al., 2025; Huang and Yang, 2025).
@@
A complementary line of work uses interactive theorem provers, where LLMs generate proofs in formal languages (e.g., Lean, Isabelle) whose correctness can be automatically checked. Key benchmarks include miniF2F (Zheng et al., 2021), FIMO (Liu et al., 2023), PutnamBench (Tsoukalas et al., 2024), which formalize competition problems, and LeanWorkbook (Ying et al., 2024), a broad Lean corpus. While this approach is promising, our work focuses on informal proofs for two reasons. First, natural language remains the primary medium for mathematical communication in education and research. Second, the task of automatically translating informal proofs into a formal language, i.e., autoformalization, is itself an exceptionally challenging research problem (Gao et al., 2025; Liu et al., 2025a).
Summary.
While prior work has established the importance of proof-based evaluation and explored the “LLM-as-a-judge” paradigm, a critical gap remains. No existing work has conducted a systematic, empirical study of the evaluator design space for mathematical proofs, nor provided a scalable methodology for creating the fine-grained, rubric-driven annotations necessary for such a study. The key factors that determine evaluator robustness remain largely uninvestigated.
@@ 6 Conclusion
We tackle the challenge of reliably evaluating natural-language mathematical proofs by introducing ProofBench, the first comprehensive, fine-grained proof grading dataset spanning multiple contests and years. Using ProofBench, we systematically explore the evaluator design space and show that strong backbones, problem-specific marking schemes, and ensembling are important, resulting in ProofGrader, our best-performing evaluator. Finally, we demonstrate practical utility by showing that using ProofGrader for best-of-$n$ selection closely tracks oracle performance.
Limitations and Future Work.
@@
While our work advances reliable evaluation of natural language math proofs, several limitations remain. First, our scope, i.e., olympiad-style proofs, does not yet cover research-level problems, or educational settings; extending to these regimes is an important next step. Second, evaluator quality can likely be improved through better prompting and workflow designs. Our dataset and analyses are intended as a foundation for such work, and we encourage future work to continue improving evaluators using ProofBench as a testbed. Our evaluation focuses exclusively on mathematical correctness. Other crucial aspects such as readability, clarity, and elegance, are not currently assessed, and we leave the development of metrics for these qualities as future work. Third, our strongest evaluators currently rely on closed models. Although our design is model-agnostic, further improving fully open-source evaluators to match those based on closed models is an important direction. While our best-of-$n$ experiments use the evaluator as a selection signal, a first step toward training-time integration, we have not yet trained with it; incorporating it into proof generator training is a promising future direction.
Acknowledgements
We thank Robin Said Sharif and Bradley Louie for their invaluable contributions to the ProofBench dataset. We also thank Jingxuan He, Zhaoyu Li, and Tony Lian for valuable discussions during the early stages of this work, as well as Prasann Singhal and Aviral Kumar for their feedback on an earlier draft of the paper.
This work was partially supported by the Google BAIR Commons Program (2025–2026) and the Gemini Academic Program Award. Sky Computing Lab is supported by gifts from Accenture, Amazon, AMD, Anyscale, Broadcom, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, NVIDIA, Samsung SDS, and SAP.
Ethics Statement
Our work adheres to the ICLR Code of Ethics. The primary ethical considerations for this research involve the creation and use of our dataset, ProofBench, and the potential for biases in our LLM-based evaluator. The dataset was constructed using problems from publicly accessible mathematics competitions. The expert annotations, which form the core of our benchmark, were provided by qualified individuals who were fairly compensated for their time and expertise.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research. To this end, we will publicly release all code, data, and supplementary materials upon publication. The ProofBench, including all problems, LLM-generated solutions, expert-annotated scores, and problem-specific marking schemes, will be made available. Our source code will include scripts to reproduce the experiments presented in this paper, including the implementation of our evaluator, ProofGraderand all prompts used.
Appendix A The Use of Large Language Models (LLMs)
In preparing this submission, we used LLMs as a general-purpose writing assistant. The human authors were responsible for the entire research process, including the initial ideation, experimental design, implementation, and data analysis. During writing, authors first creating a complete draft of each section, which included all core ideas and technical details. The LLM was then used to revise this draft for clarity, conciseness, and grammatical correctness. Its role was strictly limited to improving the language and flow of the text. All scientific contributions and claims presented in this paper originate from the authors, who take full responsibility for the final content
Appendix B Problems Information
@@
Contest Description Source
APMO The Asia Pacific Mathematical Olympiad is a regional proof-based contest for high school students across Asia–Pacific. It features five challenging problems solved over a fixed time window, proctored locally. apmo-official.org
EGMO The European Girls’ Mathematical Olympiad is an international two-day proof contest for female students, modeled after the IMO. It promotes participation and excellence among young women in mathematics. egmo.org
IMO The International Mathematical Olympiad is the premier global high school math competition. Over two days, students tackle six proof problems that test creativity, rigor, and depth. imo-official.org for problems and solutions from 2022 to 2025 and IMO 2025 problems. We use Evan Chen’s Notes for IMO 2025 solutions.
PUTNAM The William Lowell Putnam Mathematical Competition is a highly challenging proof exam for undergraduates in the U.S. and Canada. It consists of 12 problems emphasizing ingenuity and precise argumentation. maa.org/putnam
TST Team Selection Tests are national-level olympiad exams used by USA to select their IMO teams. Problems mirror international difficulty and assess readiness for global competition. Evan Chen’s Notes
USAMO The United States of America Mathematical Olympiad is an invitational proof contest following AMC/AIME qualification. It identifies top U.S. high school problem solvers and feeds into further training and team selection. Evan Chen’s Notes
@@
Table 8: Problem counts by contest and year.
Appendix C Annotation Pipeline Details
C.1 Marking Scheme Generation: Evaluation Process
We began by evaluating LLM-generated marking schemes with the objective of selecting an optimal configuration for reliable marking scheme generation. Two factors were considered: the backbone LLM and the prompting strategy (instructions).
Evaluation Protocol.
Throughout the evaluation process, two domain experts independently graded all marking schemes on a 0–3 scale using the rubric defined below. The experts compared two marking schemes at a time and rated both (the evaluation interface is shown in Figure 5).
Scoring Guidelines:
  • 0: The rubric is completely incorrect.
  • 1: The rubric seems reasonable, but the distinctions between scores and point allocations are not well-defined.
  • 2: The rubric is mostly correct, though details could be improved (e.g., inclusion of corner cases).
  • 3: The rubric is excellent—on par with those written by expert human graders.
We had the following experimental configuration:
Models: We compared two state-of-the-art models: OpenAI o3 and Gemini-2.5-Pro.
Prompts: Initial prompts were designed based on guidance from Evan Chen. [^3] https://web.evanchen.cc/static/usemo/captain-guidance-usemo.pdf Chen’s framework suggests organizing marking schemes into three sections: (1) a list of award points (indicating what progress merits credit), (2) a list of zero-credit points (for trivial observations), and (3) a list of deductions. The prompts were also designed to enforce consistent output formatting.
We compared zero-shot and few-shot prompting strategies. For few-shot prompting, examples were selected from Evan Chen’s rubrics for the United States Emergent Mathematicians Olympiad (USEMO), [^4] https://web.evanchen.cc/usemo.html as marking schemes from other competitions are not publicly available.
Stage 1 (Model Selection): We evaluated 18 problems, yielding 36 marking schemes (one per model). Zero-shot prompting was used for all generations. The two annotators differed by more than 1 point in only three cases. Overall, they preferred Gemini-2.5-Pro. Based on annotator feedback, we revised the prompts accordingly.
Stage 2 (Prompt Selection): We evaluated 36 problems, comparing zero-shot and few-shot prompting strategies. Both experts rated zero-shot prompting slightly better. Specifically, Expert 1 rated zero-shot as better in 13/36 cases, worse in 10/36 cases, and equal in 13/36 cases; Expert 2 rated zero-shot as better in 17/36 cases, worse in 8/36 cases, and equal in 11/36 cases. Feedback from Stage 2 was used to further refine the prompt design.
Average Rubric Quality: Of the 36 rubrics generated in Stage 2, 35 were rated 2 or 3 by both annotators, indicating high overall quality. Note that we did not generate or evaluate rubrics for all available problems.
Final Configuration.
The final configuration was selected based on both quantitative annotator scores and qualitative feedback. While no single setup produced perfect rubrics across all problems, we chose the configuration with the best overall performance: Gemini-2.5-Pro with zero-shot prompting.
Annotator Feedback.
We include representative feedback from the annotators below to provide insight into the rubric quality and areas for improvement:
Expert Annotator Comments General observations: Several problems included deductions for minor calculation errors, despite the prompt explicitly instructing against this practice. Due to the frequency of this issue, we did not comment on every occurrence. Problems repeated from previous batches retained preloaded feedback from earlier evaluations. Clearing this feedback between batches would prevent confusion about which problems have been completed versus those requiring re-evaluation. Many problems awarded partial credit merely for stating a lemma, definition, or formula without requiring proof or demonstration of its significance. This was the most significant issue to address in future iterations. Partial credit allocation typically followed the official solution path rather than recognizing alternative approaches that yield weaker (but still meaningful) results. For example, in USEMO-2024-P1, the AI-generated rubrics allocated partial credit across steps in the official solution, whereas the official marking scheme (by Evan Chen) awarded partial credit based on the strength of justified bounds. This discrepancy arises because the model cannot anticipate alternative solution strategies without explicit examples. A prompt modification could encourage such flexibility—for instance: “prove $X$ bound [max 4] (step 1 [additive 1], step 2 [additive 3]) (proving weaker bound $Y$ earns 2 points partial).” The “additive up to 7” scoring feature in the generated schemes does not align well with available contest rubrics. Whether this is acceptable depends on the overall objectives of the evaluation system.
C.2 Evaluation of Model-Generated Proofs
Pilot Calibration.